OpenStudy (anonymous):

How to expand log2(x^3/x-3) ?

OpenStudy (anonymous):

\[\log_{2} \frac{ x ^{3} }{ x-3 }\]

OpenStudy (anonymous):

log(a/b) = log(a) - log(b) \[\log_{2}\frac{ x^3 }{ x-3 } = \log_{2}(x^3) - \log_{2}(x - 3)\] log(a^b) = b*log(a)\[\log_{2}(x^3) - \log_{2}(x - 3)\ = 3\log_{2}(x) - \log_{2}(x-3)\]